# A Comprehensive Review of Few-shot Action Recognition

### v1

@@ VI-A More Realistic Dataset
Current research on few-shot action recognition has primarily focused on a few standard datasets (i.e., UCF [72], HMDB [73], Kinetics [8], Something to Something [77] and EPIC-Kitchens [78]), typically comprising predefined categories of actions within specific contexts. However, real-world action data is often characterized by increased complexity, diversity and noise, surpassing the coverage provided by existing datasets. Addressing this gap requires efforts toward constructing datasets with broader categories and finer-grained actions. For example, commonplace actions like “walking” can be further categorized into variations such as “walking slowly”, “walking fast” and “walking at a steady pace”. Furthermore, gathering action data across diverse scenes, including indoor and outdoor environments, various weather conditions, and different lighting conditions, is crucial for enhancing model robustness. Integrating real-world noise is also essential. Introducing background interference and noise, such as non-target actions, occlusion, camera shake, etc., effectively simulates challenges encountered in practical application environments. These efforts collectively contribute to creating more comprehensive datasets and facilitate the development of robust few-shot action recognition models.
VI-B Modality Extension
@@
Early few-shot action recognition research primarily utilizes RGB videos due to their widespread availability and accessibility. As discussed in Section V, several approaches have been developed to leverage skeleton, audio, depth, and optical flow modalities, as well as multimodal combinations, to address the FSAR task. These additional modalities can offer unique advantages for FSAR, enabling more accurate and robust recognition in various application scenarios. However, the increasing demand for recognizing daily life activities necessitates the use of additional modalities such as infrared sequences, point clouds, event streams, acceleration, radar, and WiFi. For instance, infrared sequences can capture clear action information in low-light conditions, and point clouds provide precise three-dimensional spatial structures. Future research should focus on developing effective methods for extracting features from these unexplored modalities and designing appropriate matching techniques to enhance FSAR performance. Moreover, the world is inherently multimodal, making it crucial to explore multimodal few-shot action recognition. However, the varying levels of noise and quality across different modalities complicate the integration process. Future methods need to effectively align and synchronize information across these modalities to enable realistic applications.
VI-C Adoption of Large Model
@@
Large language models (e.g., GPT-4, BERT, etc.) and multimodal large models (e.g., CLIP, DALL-E) have demonstrated remarkable performance in areas such as natural language processing and image generation. These models have been extensively pre-trained to simulate the abundant prior knowledge that humans possess before learning novel concepts with few-shot samples. This presents a valuable research opportunity to achieve better few-shot action recognition. Additionally, the representation learning capabilities of large models allow them to capture higher-level features and abstract representations of actions. In few-shot learning, leveraging pre-trained large models has the potential to facilitate faster adaptation to novel action categories and improve generalization in few-shot action recognition. Moreover, the ability of multimodal large models to effectively process and integrate information in various modalities has the potential to significantly enhance the accuracy and robustness of FSAR models, especially in complex environments with diverse action variations. Current applications [93, 94, 186] mainly aim to directly leverage semantic prior knowledge learned by large models to obtain more generalizable representations for videos. Applying large models to few-shot action recognition still holds significant promise and presents under-explored challenges.
@@
There are several issues to consider in the process of using large models. Firstly, the fine-tuning and deployment for large models are challenging. Consequently, effective model adaptation strategies are necessary to reduce computational overhead and memory requirements. Leveraging techniques such as adaptors, network compression, and dynamic architecture tuning can enhance efficiency without compromising performance. Additionally, because current large models are not pre-trained on the target domain or even on video data, addressing dataset biases and domain shifts is critical. Future research should focus on source-free domain generalization methods tailored to ensure robust performance across diverse real-world scenarios. It is also essential to prevent catastrophic forgetting to maintain generalization ability. Although large visual-language models have shown outstanding results in multiple multimodal tasks, utilizing them to effectively process multimodal video data in FSAR remains a challenge. In addition to visual and semantic information, many other modalities need to be considered in FSAR, such as audio and sensor signals. The variations across different modalities make it very complex to adjust and synchronize information across these different modalities. This encourages research to develop large models that can integrate diverse modalities for FSAR.
VII Conclusion

### v2

@@ 6.1 More Realistic Dataset
Current research on few-shot action recognition has primarily focused on a few standard datasets (i.e., UCF [125], HMDB [63], Kinetics [15], Something to Something [38] and EPIC-Kitchens [20]), typically comprising predefined categories of actions within specific contexts. However, real-world action data is often characterized by increased complexity, diversity, and noise, surpassing the coverage provided by existing datasets. Addressing this gap requires efforts toward constructing datasets with broader categories and finer-grained actions. For example, commonplace actions like “walking” can be further categorized into variations such as “walking slowly”, “walking fast” and “walking at a steady pace”. Furthermore, gathering action data across diverse scenes, including indoor and outdoor environments, various weather conditions, and different lighting conditions, is crucial for enhancing model robustness. Integrating real-world noise is also essential. Introducing background interference and noise, such as non-target actions, occlusion, camera shake, etc., effectively simulates challenges encountered in practical application environments. These efforts collectively contribute to creating more comprehensive datasets and facilitate the development of robust few-shot action recognition models.
6.2 More Realistic Settings
Currently, FSAR research predominantly follows the traditional N-way K-shot FSL setting. However, as discussed in Sec 3.1.2, there is growing interest in more general FSL settings. Therefore, it is imperative to expand FSAR into more realistic settings in future research.
Besides traditional few-shot learning tasks, significant future research avenues encompass generalized FSAR, open-set FSAR, open-world FSAR, and so on. Generalized FSAR requires the model not only to classify video samples from novel action categories accurately but also from seen categories [174]. With the advancement of large models, pre-trained video models are now capable of covering an increasing variety of action categories. This necessitates that the model rapidly generalizes to novel categories based on only a few videos, while not conflating them with a large number of seen categories. Open-set FSAR requires the model to recognize few-shot categories and reject samples that do not belong to either seen or few-shot categories. Differences between similar action classes may be subtle and temporal-dependent, thus heightening the intricacy of model recognition. This presents new challenges for the model to accurately comprehend action concepts. Additionally, open-world FSAR necessitates precise identification of actions in query videos in settings marked by scant training data and unreliable labels. Robustness is crucial for the model, which needs to accurately comprehend action concepts and remain unaffected by poor data quality. This is particularly significant in the present era of pre-training models, as the quantity and quality of samples utilized for pre-training are not consistently assured.
Inspired by the open-vocabulary action recognition task, an important research scenario for FSAR encompasses the recognition of not only the novel class (represented by support samples) to which the query sample belongs but also the specific action label. This is a highly promising and challenging task, demanding that the model accurately learn novel class concepts from few-shot samples but also exhibit strong semantic knowledge generalization capabilities. Currently, there is limited study on these FSAR settings that hold practical significance, thereby presenting another important research direction with promising applications. These emerging research directions will expand the applicability of FSAR, facilitating its effective implementation across a broader spectrum of real-world scenarios.
6.3 Modality Extension
@@
Early few-shot action recognition research primarily utilizes RGB videos due to their widespread availability and accessibility. As discussed in Section 5, several approaches have been developed to leverage skeleton, audio, depth, and optical flow modalities, as well as multimodal combinations, to address the FSAR task. These additional modalities can offer unique advantages for FSAR, enabling more accurate and robust recognition in various application scenarios. However, the increasing demand for recognizing daily life activities necessitates the use of additional modalities such as infrared sequences, point clouds, event streams, acceleration, radar, and WiFi. For instance, infrared sequences can capture clear action information in low-light conditions, and point clouds provide precise three-dimensional spatial structures. Future research should focus on developing effective methods for extracting features from these unexplored modalities and designing appropriate matching techniques to enhance FSAR performance. Moreover, the world is inherently multimodal, making it crucial to explore multimodal few-shot action recognition. However, the varying levels of noise and quality across different modalities complicate the integration process. Future methods need to effectively align and synchronize information across these modalities to enable realistic applications.
6.4 Adoption of Large Language Model
@@
Large language models (e.g., GPT-4 [1], LLAVA [80], Deepseek-R1 [39], etc.) and multimodal large models (e.g., LLAMA [136], Qwen-VL [154]) have demonstrated remarkable performance in areas such as natural language processing and image generation. These models have been extensively pre-trained to simulate the abundant prior knowledge that humans possess before learning novel concepts with few-shot samples. This presents a valuable research opportunity to achieve better few-shot action recognition. Additionally, the representation learning capabilities of large models allow them to capture higher-level features and abstract representations of actions. In few-shot learning, leveraging pre-trained large models has the potential to facilitate faster adaptation to novel action categories and improve generalization in few-shot action recognition. Moreover, the ability of multimodal large models to effectively process and integrate information in various modalities has the potential to significantly enhance the accuracy and robustness of FSAR models, especially in complex environments with diverse action variations. Current applications [159, 99, 101] mainly aim to directly leverage semantic prior knowledge learned by the language-image pretrained model to obtain more generalizable representations for videos. Applying large models to few-shot action recognition still holds significant promise and presents under-explored challenges.
@@
There are several issues to consider in the process of using large models. Firstly, the fine-tuning and deployment for large models are challenging. Consequently, effective model adaptation strategies are necessary to reduce computational overhead and memory requirements. Leveraging techniques such as adaptors, network compression, and dynamic architecture tuning can enhance efficiency without compromising performance. Additionally, because current large models are not pre-trained on the target domain or even on video data, addressing dataset biases and domain shifts is critical. Future research should focus on source-free domain generalization methods tailored to ensure robust performance across diverse real-world scenarios. It is also essential to prevent catastrophic forgetting to maintain generalization ability. Moreover, In-Context Learning (ICL) is a popular algorithm for few-shot learning especially for LLMs, still has insufficient practical application in the context of FSAR. Given the potential computational challenges that large models may encounter when processing several video data together, the adaptation of ICL methods to such models presents a challenging problem. Although large visual-language models have shown outstanding results in multiple multimodal tasks, utilizing them to effectively process multimodal video data in FSAR remains a challenge.
Recent efforts integrating Vision-Language Models (VLMs) into FSAR have predominantly relied on the CLIP model [159, 19, 59]. However, there is a lack of endeavors in FSAR utilizing video large models [190] that perform remarkable multimodal understanding and reasoning abilities. (i) The FSAR task involves the processing of video data, which typically incurs significant computational overhead. This challenge is particularly pronounced when employing VLMs. (ii) Given the continual emergence of novel action categories, the adaptation to these changes represents a crucial challenge, despite the rich common-sense knowledge possessed by VLM. (iii) Furthermore, in the context of fine-grained FSAR, it is notably challenging for Vision-and-Language Models (VLM) to capture subtle temporal differences in certain actions. In the future, it may be necessary to consider the utilization of Video Large Models from the perspective of balancing computational complexity and performance, fine-grained action recognition, and interpretability of large model decisions, as well as the design of in-context learning algorithms.
In addition to visual and semantic information, many other modalities need to be considered in FSAR, such as audio and sensor signals. The variations across different modalities make it very complex to adjust and synchronize information across these different modalities. This encourages research to develop large models that can integrate diverse modalities for FSAR.
7 Conclusion
@@
Data Availability: This survey introduces commonly used datasets for FSAR, summarized in Section 3.3. These publicly available datasets include HMDB (https://serre-lab.clps.brown.edu/resource/hmdb-a-largehuman-motion-database), UCF101 (https://www.crcv.ucf.edu/data/UCF101.php), Kinetics (https://github.com/cvdfoundation/kinetics-dataset), SSv2 (https://20bn.com/datasets/something-something), and EPIC-Kitchens (https://epic-kitchens.github.io/2021).